Chi-Square Testing 

300 employees of a company were selected at random and asked whether they were in favour of a scheme to introduce flexible working hours. The following table shows the opinion and the departments of the employees.


              OPINION          

Department In favour Uncertain Against

Production 89 42 9

Sales 53 36 11

Administration 38 12 10

 

Test whether there is evidence of a significant association between opinion and department.


1
Expert's answer
2021-08-17T17:55:56-0400

"H_0:" Opinion on flexible working hours and department are independent

"H_a:" Opinion on flexible working hours and department are dependent

"\chi^2=\sum_{\forall ij}\frac{(Oij-Eij)^2}{Eij}"

"Eij=\frac{row total(ith)\times Column total (jth)}{N}"

The table below shows the computations of Eij and chi-square




"\chi^2=\sum_{\forall ij}\frac{(Oij-Eij)^2}{Eij}"

"=0.2976+0+1.7857+0.81667+1.2+0.1+0.1111+2+2.6667"

"=8.97778"

Let "\alpha=0.05"

"df=(n-1)\times(m-1)"

"=(3-1)\times(3-1)"

"=4"

"CV=\chi^2_{4,0.05}=9.49"

"P-value=p(>9.97778)=0.06166"

Since the test statistic 8.97778 is less than the critical value 9.49, we fail to reject the null hypothesis. Similarly, since the p- value 0.06166 is greater than 0.05, we fail to reject the null hypothesis. Opinion on flexible working hours are independent of department. There is no association between opinion and department.